OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is greater than 7 and 1 die shows a 2? A 6 X 6 table of dice outcomes will help you to answer this question. Write fractions using the slash ( / ) key. Reduce fractions to their lowest terms.

OpenStudy (anonymous):

1/6 * 2/6 = 2/36

OpenStudy (anonymous):

2/36= 1/18

OpenStudy (anonymous):

correct, gratzi